5.16 What is the present value of a $700 perpetuity if the interest rate is 9%?...

5.16

What is the present value of a $700 perpetuity if the interest rate is 9%? If interest rates doubled to 18%, what would its present value be? Round your answers to the nearest cent.

Present value at 9%: $  

Present value at 18%: $  
Answer #1

Present value of perpetuity=Annual cash flows/interest rate

a.Present value=700/0.09=$7777.78(Approx).

b.Present value=700/0.18=$3888.89(Approx).